OpenStudy (anonymous):

Find the limit, if it exist, lim x--> 0 (1/2+x-1/2) /x

OpenStudy (anonymous):

\[\lim_{x\to0}\frac{\frac{1}{2}+x-\frac{1}{2}}{x}=\lim_{x\to0}\frac{x}{x}=\lim_{x\to0}1=\cdots\]